Suppose a recurrence relation


an=an−1+20an−2

where a1=9 and a2=189


can be represented in explicit formula, either as:


Formula 1:

an=pxn+qnxn

                      or  


Formula 2:

an=pxn+qyn

 


where 

x

and

y

are roots of the characteristic equation.


**If the explicit formula is in the form of Formula 2, consider p > q.


 


Determine p and q

Answer:


1
Expert's answer
2021-12-06T16:27:07-0500

Let us solve the characteristic equation of the recurrence relation "a_n=a_{n\u22121}\u221220a_{n\u22122}," which is equivalent to "a_n-a_{n\u22121}-20a_{n\u22122}=0." 

It follows that the characteristic equation "x^2-x-20=0" is equivalent to "(x+4)(x-5)=0," and hence has the roots "x_1=-4, x_2=5." 

It follows that the solution of the recurrence equation is 

"a_n=p\cdot (-4)^n+q\cdot (5)^n."

 Since "a_1=9" and "a_2=189," we conclude that


"9=a_1=-4p+5q""189=a_2=16p+25q"


"9=-4p+5q""36+189=-16p+16p+20q+25q"


"9=-4p+5q""225=45q"

Therefore, "q=5" and "p=4."



"a_n=4(-4)^n+5(5)^n."